Consider the R 2 − R function f defined by f(x,y) = x- 2y.

Prove from first principle that

lim f (x, y) = 0.

(x,y)→(2,1)



1
Expert's answer
2021-05-07T14:18:30-0400

Given function is-

"f(x,y)=x-2y"


Let "\forall \in >0" be given


Now consider ,

    "|f(x,y)-f(2,1)|\n\n =|x-2y-0|=|x-2y|"


"\Rightarrow |x-2-2(y-1)|



\Rightarrow |x-2-2(y-1)|\le |x-2|+2+y-1|"


If we choose "|x-2|<\in, |y-1|<\in"


then "|x-2'-2(y-1)|<\in+2\in=3\in"


   "|x-2-2(y-1)|<3\in=\in"


So, "\in>0, \exist \in >0" such that


for "|x-2|<\in, |y-1|<\in \Rightarrow f(x,y)-0|<\in"


i.e. "lim_{(x,y)\to (2,1)}f(x,y)=0"